What is the Answer and Easiest way to Solve these Equations:
x(x+8)(x+7)=0 , y(y-1)(2y+4)=0 , (n-7)Squared(5n+20)=0 , (c-2)(2c+8)(5c+10)=0

Answer by stanbon(75887)

What is the Answer and Easiest way to Solve these Equations:
If any one of those factors is zero the product is zero.
------
x(x+8)(x+7)=0
x = 0 or x = -8 or x = -7
--------------------------------
y(y-1)(2y+4)=0
y = 0 or y = -2
-------------------------
(n-7)Squared(5n+20)=0
n = 7 or n = -4
------------------------
(c-2)(2c+8)(5c+10)=0
c = 2 or c = -4 or c = -2
=============================
Cheers,
Stan H.
==================